Mir Status Report 20 November 1997

NASA MIR Program / Mir 24 / NASA 6 Status
Date: Thursday, November 20, 1997 // MD 107/56

Mir Station System/Subsystem Activities

Today's major activities centered on preparation and execution of the Cooperative Solar Array experiment. The test consisted of disconnecting, measuring, and connecting multiple solar cells on the Cooperative Solar array (Kvant -Z axis) and the disconnection and re-connection of the Russian Solar array (Kvant +Z axis - installed during the November EVA). Although the Base Block arrays remained connected during the entire test, they were shadowed by Kvant arrays due to the Beta angle and attitude of the experiment. As a precaution, Dave was told to power off IPCG and Glovebox during the comm pass at 18:15 DMT. Additionally, Dave was told to power off PUP A and PUP B for the night passes from 19:00 until 00:15.

At the most critical time of the test, the power margin turned out to be about 20 amps. Mr. Blagov expressed his thanks to the Phase I Program for agreeing to power down some of the science equipment and to cycle the PUP's power.

The only other item of interest we are tracking is the temp in the Base Block. The crew called down temps of 36 degrees (18:16 DMT) and 30 degrees (22:30 DMT). The ground team is attributing the high temperatures to off nominal performance of the Base Block cooling loop KOB-1. The ground is also assessing powering the Base Block air conditioner to make the sleep period more comfortable. With the higher temperatures, the Kvant Electron remained off. The crew reported that they used three oxygen canisters to maintain proper ppO2 levels.

At the 00:11 comm pass on Nov 20/21: The cosmonauts were finishing the the reconnections of the all of the cables and closing out of the panels.

The ground instructed them to turn on both condensate removal pumps. They did and reported that they were operational, even the #2 pump which surprised them. Apparently, there were problems with this pump earlier. Later in the pass, the ground reported that the condensate pumps were not operational per the tm. The ground told the crew this can wait 'til tomorrow.

The ground also instructed the cosmonauts to burn one more candle which gives a total of 3 O2 candles today.

The crew reported that the solar array tracking control was reinitiated. Later in the pass, the crew reported that the solar array drives were nominal

Before the comm pass the Power flight controllers were asking for last minute changes to the attitude, which the SFD took objection to, but gave in. The Mir was in LVLH attitude and the power flight controller wanted to place the station into solar inertial. The fact that the Mir was in LVLH was surprise to us, so Joel has added this to his list of topics for discussion with Yuri Antoshechkin.

At the Post-Comm Pass SFD Conference:

- MCS performing nominal

Base Block - nominal
Kvant 2 Module - Flight Controller reported an anomaly, but this was based on
the previous rev telemetry. The telemetry for this pass was not received for
the Kvant 2.

-C&T: Svet and Symbol - nominal

Mir Payload Activities
MG:FAC:I:39 BEM/BSTC-M and BTR Facility Maintenance
FB:HSR:I:23 CHAPAT Experiment Computer Monitoring
MG:BT:I:16 BIO3D Red Cell Line Fixation and Stowage
MG:BT:I:90 BIO3D Blue Cell Line Sample Transfer, Fixation and Stowage of
Fixed Samples
MG:BT:I:91 BIO3D Green & Gold Cell Lines Fixation and Stowage
SMP:ENV:I:48 Activate SSAS
SMP:ENV:I:49 GRAB Air Sampling
SMP:ENV:I:56 MO-14 Formaldehyde Monitoring
SMP:MO:I:44 MO-11 TEPC Display Call Down
HLS:MET:I:92 Night Headband Postsleep*
HLS:MET:I:101 Night Headband Presleep*
EVA FDF Study (2 Hours)
*Cosmonaut Activities

Experiment Operations
BIO-3D - There are currently no hardware anomalies with either the BSTC or
the BTR.

TEF's - Were NOT de-energized.

SMP/Medical -- Dave performed environmental monitoring today with the start of Formaldehyde and SSAS and the completion of a grab air sample. A TEPC calldown was also scheduled to be performed. The temperature on board MIR has risen to 30 C and is being monitored. The plan is to run the A/C tonight after the completion of the solar array test.

Environmental Parameters (previous 24 hour period):
Parameter Min. Max. Nominal Range
P (mmHg) BB 668 670 660 - 850
ppO2 (mmHg) 151 152 140 - 200
ppCO2 (mmHg) 3.4 3.8 up to 8.0
ppH20 (mmHg) 14.0 16.2 8 - 14
Humidity (%) 45 51 30 - 75
Temp (xC) BB 29.5 30.3 18 - 28

Priroda Temperature (deg C) at BV1 Fan: 27.3

  • submit to reddit


Support SpaceRef, NASA Watch and the Astrobiology Web on Patreon.

Join our mailing list

Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on November 20, 1997 9:00 PM.

Mir Status Report 19 November 1997 was the previous entry in this blog.

Read Dan Goldin's lips: Better-Cheaper-Faster is Better-Cheaper-Faster is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.